OpenStudy (anonymous):

lowest term of 10/14

OpenStudy (shamil98):

Both can be reduced by two.

OpenStudy (shadowlegendx):

Divide each number by 2 You should get 5/7

OpenStudy (australopithecus):

A fraction will not change in value as long as you multiple the top and the bottom, or divide the top and the bottom by the same thing, for instance: \[\frac{10}{2} = \frac{10/2}{2/2} = \frac{5}{1} = 5\]

OpenStudy (australopithecus):

This is an important rule to keep in mind when doing algebra

OpenStudy (zale101):

@chastin

OpenStudy (mathmale):

Chastin: Mind explaining what you mean by "lowest term of 10/14"? My mind immediately interprets that as "reduce 10/14 to lowest terms."